OpenStudy (anonymous):

Complete the reaction. Write your balanced equation two different ways C6H14 + O2

OpenStudy (asib1214):

when it says two different ways it means that you need to write the equation of both complete and incomplete reaction of the compound. You know that when a compound is burned under very limited amount of oxygen, it makes an incomplete combustion. The following are the products of an incomplete combustion reaction...Carbon, Carbon Monoxide, Carbon Dioxide, H2O and energy. but when it's a complete reaction, the products are Carbon Dioxide and water. That's the reason the question is asking you to balance the equation in two ways, in a complete reaction and second in incomplete reaction

OpenStudy (asib1214):

wanna try first???

OpenStudy (anonymous):

2 C6H14 + 19 O2 --> 12 CO2 + 14 H2O is the complete reaction?

OpenStudy (anonymous):

I have no idea what the incomplete reaction is.. I don't think that was explained in class.. or I fell asleep.

OpenStudy (asib1214):

is the chemical equation above is balanced then it's an equation of a complete reaction...

OpenStudy (asib1214):

now what about the incomplete reaction???? wanna try

OpenStudy (anonymous):

yep. i know that one is the complete one. i think O2 has to be alone for incomplete..? sorry, no idea

OpenStudy (asib1214):

ok i'm going to write down the chemical equation but you need to balance it own your own.

OpenStudy (anonymous):

no oxygen and hydrocarbon have to be fuels

OpenStudy (anonymous):

okay I'm pretty good at balancing

OpenStudy (anonymous):

i know the left side has to have the same C6H14 + O2

OpenStudy (asib1214):

when there is less oxygen available for a combustion reaction, the hydrocarbon breaks down into very dangerous substances as i've mentioned above. C + CO + CO2 + H2O + energy.

OpenStudy (anonymous):

ah okay

OpenStudy (asib1214):

now try to balance this.....C6H14 + O2 = C + CO + CO2 + H2O

OpenStudy (anonymous):

alright, hold on

OpenStudy (anonymous):

hmm i still get: 2 C6H14 + 19 O2 = 12 CO2 + 14 H2O

OpenStudy (anonymous):

but that can't be right...

OpenStudy (anonymous):

the general equation does have to look like : C6H14 + 02 → C02 + H20

OpenStudy (asib1214):

wait let me balance the equation for you, but try to do the next own your own....

OpenStudy (anonymous):

2 C6H14 + 19 O2 = 12 CO2 + 14 H2O i keep getting this

OpenStudy (asib1214):

C6H14 + 02 → C02 + H20 this equation represents a complete reaction in which both reactants completely react with each other and produce CO2 + H2O. In this situation the oxygen was plentyful that's why it didn't break into a 4 separate compounds i mentioned above.

OpenStudy (asib1214):

i'm going to balance the incomplete combustion equation for you.

OpenStudy (anonymous):

okay thanks!

OpenStudy (asib1214):

ok here it is.... 2C6H14 + 13O2 = 4C + 4CO + 4CO2 + 14H2O C=12 C= 12 H=28 H= 28 O=26 O= 26

OpenStudy (asib1214):

that's the second way of balancing the equation....

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

i understand now!